OpenStudy (anonymous):

Find the altitude of an equilateral triangle, if each side of the triangle has a length of 14 meters.

OpenStudy (lucaz):

the altitude of an equilateral triangle with sides of 1 unit is sqrt(3)/2, proportionaly 14*1=14 and 14*sqrt(3)/2*=7sqrt(3)

OpenStudy (anonymous):

There is another way. In a equilateral triangle, each angle is 60 deg. Call h the height, and a the side = 14. Trig table gives sin 60 = V3/2 sin 60 = h/a = h/14 -> h = 14*(V3/2) = 7V3